OpenStudy (anonymous):

the reciprical of 4 less than a # is twice the reciprical of the # itself. what is the #

OpenStudy (anonymous):

u dont have to type it all out

OpenStudy (anonymous):

lets say the number is x what is 4 less than the number? it is x-4 what is the reciprocal of 4 less than the number? it is 1/(x-4) what is the reciprocal of the number itself? it is 1/x what is 2 times the reciprocal of the number itself? it is 2*1/x = 2/x so 1/(x-4) = 2/x solve for x